x sqrt 3 -2=5x sqrt 3
I am having a mental problem and need help with this one

Answer by mananth(16946)

x sqrt 3 -2=5x sqrt 3
5XSQRT3-XSQRT3=-2
4XSQRT3=-2
X= -2/4 SQRT3
X= -1/2 SQRT3